OpenStudy (anonymous):

5. Solve, the following system of equations by linear combination: 2d + e = 8 d -e = 4

OpenStudy (anonymous):

2d +e =8 d-e=4 then d=4+e substitute 2(4+e)+e=8 8+2e+e=8 3e=0 e=0 so d=4

OpenStudy (anonymous):

2d+e=8 d-e=4 --------- d= 4

OpenStudy (anonymous):

is that how you do this?

OpenStudy (anonymous):

i kinda for got how do do combinations they dont pop up too much

OpenStudy (anonymous):

i know me too!

OpenStudy (anonymous):

The solution is (5, 2). The solution is (4, 0). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

those are the choices

OpenStudy (anonymous):

if you are saying that there are an infinite number of sol then there is a sol

OpenStudy (anonymous):

just many

OpenStudy (anonymous):

the answer no sol. would be incorrect if that was answered on a test question

OpenStudy (anonymous):

ohh ok thanks

OpenStudy (anonymous):

well gtg now sleep

OpenStudy (anonymous):

bye bye now

OpenStudy (anonymous):

night!

OpenStudy (anonymous):

:)